A boy weighing 445 N swings on a 2-m long swing. If his horizontal speed at the lowest point is 3 m/s, what total force must the ropes holding the swing be able to withstand?


1
Expert's answer
2021-05-21T10:05:22-0400

"T-mg=m\frac{v^2}{r}\to T=mg+m\frac{v^2}{r}=W+\frac{W}{g}\cdot\frac{v^2}{r}="


"=445+\frac{445}{9.8}\cdot\frac{3^2}{2}\approx650\ (N)" . Answer